The label on a stock bottle of HCl reads: Specific gravity 1.18; 36% HCl. How would you prepare 0.5dmcube of 3.0M HCl
1
Expert's answer
2021-08-11T05:22:40-0400

Molar Mass of HCL = 36.458 g/mol

Molar Mass of 36% = "\frac{36}{100}×36.458 g/mol = 13.12g/mol"

"1dm^3= 1000ml"

"0.5dm^3= \frac{1000×0.5}{1}=500ml"

"\frac{3×500}{13.12g}=114g"

To prepare 3M of HCL you use 114 g